Suppose that a firm has only one variable input, labor, and firm output is zero when labor is zero. When the firm hires 6 workers the firm produces 90 units of output. Fixed costs of production are $6 and the variable cost per unit of labor is $10. The marginal product of the seventh unit of labor is 4. Given this information, what is the marginal cost of production when the firm hires the 7th worker?
1
Expert's answer
2018-04-11T09:48:11-0400
FC=$6
VC=$10
When L=6,Q=90
〖MPL〗_7=4
〖MP〗_L=∆Q⁄∆L
4=∆Q⁄1
∆Q=4
MC=∆TC⁄∆Q
〖TC〗_6=FC+VC=6+10*6=66
〖TC〗_7=FC+VC=6+10*7=76
∆TC=76-66=10
MC=10⁄4=2.5
Answer: $2.5